Using activities, calculate the concentration Cu2+ if the overall potential of the cell is 0.572V. Show all the work and state your final result with 3 sig fig. The standard reduction potentials are: E Cu2+/Cu(s)=0.339V and ENi2+/Ni(s)=-0.236V. Assume an activity coefficient for Cu2+=0.690 and for Ni2+=0.0675
Answer #1

In this problem a cell is made up of Cu and Ni electrodes,

Standard reduction potential: Cu2+ / Cu = 0.339V, and

Ni2+/ Ni = -0.236 V , it has comparatively lower reduction potential, so it will act like an anode and have oxidation on it, with standard oxidation potential = +0.236 V, while the other electrode is acting as cathod.

The reaction involving in this problem is:

At Anode- Ni(s) \rightarrow Ni^{^+2} + 2e^{-}

At cathod - Cu^{^+2} + 2e^{-} \rightarrow Cu(s)

Overall reaction - Ni(s) + Cu^{^+2} \rightarrow Ni^{^+2} + Cu(s)

Overall EoCell=(0.236 + 0.339) V = 0.575 V

Formula of EMF or  overall potential of the cell:

E_{cell} = E^{o}_{cell} - \frac{2.303RT}{nF} log \frac{[Ni^{+2}]}{[Cu^{+2}]}

Since we have not provided with temperature of the reaction, so we are assuming reaction is taking place at room temperature, for that, \frac{2.303RT}{nF} = \frac{0.059}{n}

therefore, (Ni)a 0.059 log- (Cu2+)a Ecell= Ecell

here, n= number of electrons involving in this process = 2 & Ecell = 0.527 V

0.059 (Ni2+)a -log 2 (Cu2+)a 0.527 0.575

(Ni)a - 1.627 og (Cu2+a

(Ni)a 42.364 Cu2+a

(Ni2+)a and (Cu2+)a are the activities of Ni2+ and Cu2+ ions, respectively, and

Concentration x Activitycoef ficient Activitity

activity coefficient for Cu2+=0.690 and for Ni2+=0.0675.

Ni2+x 0.0675 Cu2+] x 0.690 = 42.364

433.054 Cu2+

Cu2+ =0.0023 x [Ni2